A solution is prepared by mixing equal volumes of 2.0 ´ 10-4 M Barium chloride (BaCl2) solution and 3.0 ´ 10-3 M sodium phosphate (Na3PO4 ) solution at 25°C. Would Ba3(PO4)2precipitate out? Ksp of Ba3(PO4)2 at 25°C is 3.0 ´ 10-23.
Answer #1

sol: the ballanced equation is

3BaCl2 + 2Na3​​​​​​PO4​​​​​​ = Ba3(PO4)2 + 6NaCl

For 3 moles of BaCl2  2 moles of Na3​​​​​​PO4 is required

So for 1 mole of BaCl2  2/3 = 0.67 moles of